Checking to see if my highlighted answers are correct????

__d__ 1. What is a standard cost?

a.   The total number of units times the budgeted amount expected

b.   Any amount that appears on a budget

c.   The total amount that appears on the budget for product costs

d.   The amount management thinks should be incurred to produce a good or service

__c__ 2. A standard cost is

a.   a cost which is paid for a group of similar products.

b.   the average cost in an industry.

c.   a predetermined cost.

d.   the historical cost of producing a product last year.

__c__ 3. The difference between a budget and a standard is that

a.   a budget expresses what costs were, while a standard expresses what costs should be.

b.   a budget expresses management's plans, while a standard reflects what actually happened.

c.   a budget expresses a total amount, while a standard expresses a unit amount.

d.   standards are excluded from the cost accounting system, whereas budgets are generally incorporated into the cost accounting system.

PART II — Direct Labor Standard (20 points)

Engines Done Right Co. is trying to establish the standard labor cost of a typical engine tune-up. The following data have been collected from time and motion studies conducted over the past month.

                     Actual time spent on the tune-up                 1.0 hour

                     Hourly wage rate                                          $16

                     Payroll taxes                                                10% of wage rate

                     Setup and downtime                                    10% of actual labor time

                     Cleanup and rest periods                            20% of actual labor time

                     Fringe benefits                                             25% of wage rate

Instructions

(a)     Determine the standard direct labor hours per tune-up

standard direct labor hours per tune-up = Actual time spent + Setup and downtime +          Cleanup and rest periods

standard direct labor hours per tune-up = 1 + 1*10% + 1*20%

standard direct labor hours per tune-up = 1.3 Hours

(b)     Determine the standard direct labor hourly rate.

         standard direct labor hourly rate = Hourly wage rate + Payroll taxes + Fringe benefits

         standard direct labor hourly rate = 16 + 10%*16 + 25%*16

         standard direct labor hourly rate = 21.60

(c)     Determine the standard direct labor cost per tune-up.

         standard direct labor cost per tune-up = standard direct labor hours per tune-up*standard     direct labor hourly rate

         standard direct labor cost per tune-up = 1.3*21.6

         standard direct labor cost per tune-up = $ 28.08

  (d)     If a tune-up took 1.5 hours at the standard hourly rate, what was the direct labor quantity variance?

        Direct labor quantity variance = (Actual Hour - Standard Hour) * Standard Rate

        Direct labor quantity variance = (1.50 - 1.30)*21.60

        Direct labor quantity variance = $ 4.32 Unfavorable
Answer #1

__d__ 1. What is a standard cost?

Ans :- d.   The amount management thinks should be incurred to produce a good or service

       because management think or determine what is the cost should be incurred to produce product

__c__ 2. A standard cost is

Ans: - c.   a predetermined cost

          because management are thinking what cost should be incurred to produce product therefore standard cost is a predetermined cost.

__c__ 3. The difference between a budget and a standard is that

Ans: c.   a budget expresses a total amount, while a standard expresses a unit amount.

yes its true or right

(a)     Determine the standard direct labor hours per tune-up

standard direct labor hours per tune-up = Actual time spent + Setup and downtime +Cleanup and rest periods

standard direct labor hours per tune-up = 1 + 1*10% + 1*20%

standard direct labor hours per tune-up = 1.3 Hours

Note :- this answer is right because total hour per tune incurred from start point to end point means from setup time to finish time

(b)     Determine the standard direct labor hourly rate.

         standard direct labor hourly rate = Hourly wage rate + Payroll taxes + Fringe benefits

         standard direct labor hourly rate = 16 + 10%*16 + 25%*16

         standard direct labor hourly rate = 21.60

Note :- this answer is right standard hours rate includes all cost related to the total product process cost like wages, taxes , transport cost and other expenses

(c)     Determine the standard direct labor cost per tune-up.

         standard direct labor cost per tune-up = standard direct labor hours per tune-up*standard     direct labor hourly rate

         standard direct labor cost per tune-up = 1.3*21.6

         standard direct labor cost per tune-up = $ 28.08

Note:- This answer is right or correct

  (d)     If a tune-up took 1.5 hours at the standard hourly rate, what was the direct labour quantity variance?

        Direct labour quantity variance = (Actual Hour - Standard Hour) * Standard Rate

        Direct labour quantity variance = (1.50 - 1.30)*21.60

        Direct labour quantity variance = $ 4.32 Unfavourable

Note :- this answer is right because actual hour is more than standard hours therefore cost of product will increase therefore it is unfavourable

all the answer are right please see my notes.